OpenStudy (moonlitfate):

Use the geometric power series along with differentiation or integration to determine a power series for the function.

OpenStudy (moonlitfate):

Geometric Power Series:\[\frac{ 1 }{ 1-x } = \sum_{0}^{\infty}x^n , |x|<1\] Function: \[f(x)= \frac{ 2 }{ (1-2x)^2 } = \frac{ d }{ dx }[\frac{ 1 }{ 1-2x }]\]

OpenStudy (freckles):

do you know the power series for 1/(1-2x)?

OpenStudy (moonlitfate):

No, but couldn't manipulate 1/(1-2x) to try and figure that out? This stuff just really confuses me.

OpenStudy (freckles):

you know that \[\frac{1}{1-u}=\sum_{n=0}^{\infty} u^n, |u|<1\] replace u with 2x

OpenStudy (moonlitfate):

Oh, wow. Okay. So I would just have to substitute 2x everywhere that I would I see an x in the original P. Series.

OpenStudy (freckles):

\[\frac{1}{1-2x}=\sum_{n=0}^{\infty} (2x)^n, |2x|<1 \\ \text{ \right but now you have }\\ f(x)=\frac{d}{dx}(\sum_{n=0}^{\infty} (2x)^n) , |2x|<1\]

OpenStudy (freckles):

which means you need to find the derivative term by term or just cheat and find the derivative of (2x)^n

OpenStudy (freckles):

with respect to x of course

OpenStudy (freckles):

lol it isn't really cheating

OpenStudy (moonlitfate):

f'(x) = 2n(x)^(n-1)

OpenStudy (moonlitfate):

Hopefully my deriving skills aren't that rusty...

OpenStudy (freckles):

well..It looks like you used chain rule but you didn't leave the inside the same

OpenStudy (freckles):

and also you know to have the sum thing around that right?

OpenStudy (moonlitfate):

Oh.. right. Okay, so it would look like this then: \[\sum_{n=0}^{\infty}2n(2x)^{n-1}\]

OpenStudy (freckles):

and one more think we would not say f'(x) equals but f(x) equals.. \[\frac{1}{1-2x}=\sum_{n=0}^{\infty} (2x)^n, |2x|<1 \\ \text{ \right but now you have }\\ f(x)=\frac{d}{dx}(\sum_{n=0}^{\infty} (2x)^n) , |2x|<1 \\ f(x)=\sum_{n=0}^{\infty}2n(2x)^{n-1}, |2x|<1\]

OpenStudy (freckles):

right

OpenStudy (freckles):

and you can solve that one inequality for x in for you want to or solve for |x|

OpenStudy (freckles):

and then you are done :)

OpenStudy (moonlitfate):

That's it? :) These problems like so ridiculous. Just makes your brain hurt, lol.

OpenStudy (moonlitfate):

*look

OpenStudy (freckles):

you can leave that little thing inside the sum as is or do this: \[2n(2x)^{n-1} \\ 2n \cdot 2^{n-1} x^{n-1} \\ 2^{n-1+1}n x^{n-1} \\ 2^{n} nx^{n-1}\] but no big deal

OpenStudy (freckles):

the way we have it above is fine though

OpenStudy (moonlitfate):

Ah, okay. :) I'll work on it, and if I get stuck I'll just send you a message, if that's okay?

OpenStudy (freckles):

alright

OpenStudy (moonlitfate):

Thank you so much.

OpenStudy (freckles):

np